Consider the general chemical equation 3A+B→2C.
If 1.55 g of A reacts with 1.55 g of B, what is the mass of C?
Express your answer with the appropriate units.

If 4.00 g of A reacts to produce 8.95 g of C, what is the mass of B?
Express your answer with the appropriate units.
1
Expert's answer
2020-11-23T06:57:14-0500

According to the law of conservation of mass;


Total mass of reactants = Total mass of product


1. Total mass of reactants = 3A + B = 1.55 + 1.55 = 3.1g


Total mass of products is therefore = 2C = 3.1g


The mass of C produced in the reaction is 3.1 g



2. Total mass of product produced = 8.95g

Mass of A = 4.00g


According to the law of conservation of mass,


mass of B = mass of mass of products - mass of A = 8.95g - 4.00g = 4.95g


The mass of B is 4.95 g